# TrackGS: Optimizing COLMAP-Free 3D Gaussian Splatting with Global Track Constraints

### v2

@@ 5.1 Experimental Setup
Table 1: Novel view synthesis results on Tanks and Temples and CO3D-V2. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are gray background.
During the training phase, we will clone new Gaussians from those associated with the track points and apply the same training strategy as the original 3DGS (including clone, split, and delete). Note that the track Gaussians still need to be preserved and constrained by a scale loss. We use Eq. 13 for training and set $\lambda_{1}=0.8$, $\lambda_{\mathrm{D-SSIM}}=0.2$, $\lambda_{t2d}=0.01$, $\lambda_{t3d}=0.01$, $\lambda_{scale}=0.01$.
Figure 5: Qualitative comparison for novel view synthesis and camera pose estimation on CO3D-V2. Benefit from the accuracy of the camera pose estimation, the rendering quality of novel view synthesis obtained by our method is higher than CF-3DGS.
5.2 Experimental Results and Analysis
Novel View Synthesis. Unlike the standard settings where the camera poses of test views are given, we need to first obtain the camera poses of the test views for rendering. Following the same protocol as CF-3DGS [6], we obtain the camera poses of the test views by minimizing the photometric error between the synthesized images and the test views using the pre-trained 3DGS model. We apply the same procedure to all baseline methods to maintain a consistent bias for a fair comparison.
We report the results on Tanks and Temples and CO3D-V2 in Tab. 1. Our method consistently outperforms NoPe-NeRF across all metrics and slightly outperforms CF-3DGS in overall performance on Tanks and Temples. On CO3D-V2, our results are better than those of all baselines. Interestingly, compared to HT-3DGS [9], our method achieves significantly higher PSNR on the CO3D-V2, which involves large camera motion, but performs worse on the Tanks and Temples, which features smooth motion trajectories. We attribute this to HT-3DGS’s reliance on video frame interpolation, which can be unstable under large camera motion. As shown in Fig. 3, the images synthesized by our method are clearer than those obtained by other methods, as evidenced by higher scores in terms of SSIM and LPIPS, as detailed in Tab. 1. As illustrated in Fig. 4 and  5, the advantages of our algorithm are well demonstrated, especially with large camera motions. Due to global joint optimization, multi-view geometric consistency is better maintained in the trained 3DGS model, leading to high-quality images.
For further comparison, we evaluated our method on the Synthetic Dataset, which features extremely complex camera motions. One result is shown in the bottom-right of Fig. 1. In this case, the camera not only moves in multiple circles around the object but also changes significantly in the vertical direction. Our synthesized image from the novel view remains clear and sharp, whereas the CF-3DGS result is blurry with obvious artifacts.
Scenes Ours HT-3DGS CF-3DGS NoPe-NeRF
@@ 5.2 Experimental Results and Analysis
Table 2: Quantitative comparison of pose accuracy on CO3D-V2. Note that, we use COLMAP poses as the ground truth. The unit of $\mathrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\mathrm{RPE}_{t}$ is scaled by 100. The best results are gray background.
Scenes Ours CF-3DGS COLMAP+3DGS
[…]
classroom 36.26 0.94 0.15 19.69 0.69 0.46 35.81 0.94 0.15
lego_c2 29.36 0.90 0.12 15.93 0.31 0.55 28.77 0.88 0.15
livingroom 33.52 0.88 0.24 16.63 0.57 0.57 32.74 0.87 0.27
bedroom 31.17 0.93 0.13 16.98 0.65 0.45 31.73 0.94 0.13
Table 3: Novel view synthesis results on our Synthetic dataset. We show the quantitative results using our method, CF-3DGS and COLMAP+3DGS. The best results are gray background.
Camera Parameter Estimation. First, we compare the camera pose estimation with baseline methods. In the comparison, our method only assumes that the camera focal length is fixed, while others additionally input the camera focal length. The estimated camera poses are analyzed by Procrusts as in CF-3DGS and compared with the ground-truth of training views. The quantitative results of camera pose estimation on CO3D-V2 datasets are summarized in Tab. 2111Due to the hierarchical training of HT-3DGS, it needs significant computational resources and causes OOM errors in our experimental setup. We report results from the paper, only covering partial scenes from CO3D-V2.. The results show that our estimated camera parameters achieve the smallest error among all methods, with the Absolute Trajectory Error (ATE) being only one-fourth of that of the second-best method. This demonstrates that our algorithm excels in scenes with complex camera motions. Compared to the baselines, the global tracking information we use eliminates accumulated errors, leading to more accurate camera pose estimation. Additionally, joint optimization enhances the stability of the estimation results.
Next, we evaluate camera parameter estimation on our Synthetic Dataset. Tab. 4 shows the estimation errors from CF-3DGS, COLMAP, and ours. Note that CF-3DGS uses the camera FoV estimated by COLMAP. We find that our estimated camera FoVs and poses are comparable to those of COLMAP, and the camera pose error is 10 times smaller than that of CF-3DGS. Fig. 6 visualizes our estimated poses in different stages. Thanks to the joint optimization based on global track and the back-propagation of the gradient of the camera parameters, our approach is able to combine these two tasks, thereby reducing the input requirements.
Algorithm Efficiency. All experiments were conducted on a single RTX 4090 GPU. On average across all scenes, the processing times for CF-3DGS and NoPe-NeRF are approximately 1.5 hours and 4 hours, respectively. HT-3DGS reports a runtime of around 4 hours on a professional-grade GPU. In contrast, our approach takes less than 1 hour and uses less than 12GB of GPU memory on average, as we optimize all parameters, including camera parameters.
@@
[…]
CF-3DGS 0.993 0.588 2.436 0.07412 0.021 1.1265 4.946 0.10795 0.029 0.425 2.104 0.07653 0.042 0.366 1.103 0.06284
COLMAP+3DGS 0.993 0.004 0.018 0.00007 0.021 0.009 0.026 0.00019 0.029 0.008 0.026 0.00014 0.042 0.009 0.035 0.00023
@@
Table 4: Quantitative comparison of parameter accuracy on our Synthetic dataset. We convert the estimated camera intrinsics focal to FoV and perform the errors of FoV with ground truth (provided by our synthetic datasets) on our method, CF-3DGS and COLMAP+3DGS. As CF-3DGS requires the camera intrinsic parameters as fixed inputs, we set them the same as COLMAP+3DGS.
5.3 Ablation Study
Effectiveness of Different Losses. We ablate each loss of the algorithm on synthetic dataset, since it has ground-truth camera parameters. Tab. 5 reports the average synthesis quality and camera parameter errors across different algorithm variants (see supplementary material for details). In order to better show the role of each loss, we construct variants of the three schemes and remove them one by one. First, we can find that any variant of our algorithm (Variant 3, 4, 5, 6) is better than the CF-3DGS method (Variant 3) in synthesis quality and absolute camera position. Second, from Variant 3, it shows that 2D track loss plays a crucial role in the entire joint optimization. When 2D track loss is not used, compared with the final method (Variant 3 vs. 6), there is a significant decrease in synthesis quality (18.18 vs. 32.58), and the camera parameter error is significantly larger (2.617 vs. 0.015). This shows that the reprojection error constrained by global consistency can significantly enhance the camera parameter estimation, thereby improving the 3DGS training effect and improving the new perspective synthesis ability. In addition, the results of Variant 4 vs. 6 show that 3D track loss can further enhance the geometric consistency of 3DGS. When 3D track loss is used, the PSNR of the novel view synthesis can be further improved by 0.2 dB, and the absolute position error of the camera can be reduced by an order of magnitude. From the experiment, we find that when scale loss is not used, Variant 5 has obvious degradation in all results of the entire scene. This is related to the fact that our method achieves explicit tracking by limiting the size of the tracked Gaussian and using it as a virtual spatial 3D point. Tab. 6 further illustrates the role of scale loss in constraining the size of the tracked Gaussian. It can be seen that after using scale loss, the centroid of the Gaussian is an order of magnitude closer to the surface than those without it.
We further analyze the role of track losses (2D&3D) and scale loss on the CO3D-V2 in Tab. 7. Additionally, we visualize the camera poses and rendered images with and without track losses in Fig. 8. Our observations indicate that global tracks are crucial for improving both novel view synthesis and pose estimation, as the tracks enforce multi-view geometric consistency during 3DGS training.
@@ 5.3 Ablation Study
Table 5: Ablation study on different losses.
classroom lego_c2 livingroom bedroom
[…]
Table 6: Effectiveness of scale loss in regulating Gaussian sizes.
scenes w.o. 2D&3D track w.o. scale Ours
[…]
34_1403_4393 28.23 0.86 0.0179 28.80 0.87 0.0110 28.68 0.88 0.0088
46_2587_7531 30.10 0.90 0.0110 31.64 0.91 0.0006 31.83 0.92 0.0005
106_12648_23157 19.21 0.47 0.0026 21.97 0.63 0.0014 26.18 0.83 0.0009
110_13051_23361 25.13 0.73 0.0208 33.04 0.94 0.0002 33.44 0.94 0.0002
245_26182_52130 30.38 0.80 0.0099 34.00 0.93 0.0001 33.82 0.93 0.0001
407_54965_106262 27.35 0.83 0.0180 29.20 0.86 0.0090 28.73 0.86 0.0100
415_57112_110099 23.79 0.62 0.0041 23.49 0.61 0.0014 30.37 0.88 0.0010
429_60388_117059 24.57 0.62 0.0108 24.91 0.65 0.0096 25.70 0.70 0.0092
mean 26.10 0.73 0.0119 28.38 0.80 0.0042 29.84 0.87 0.0038
Table 7: Effectiveness of track losses and scale loss. Performance on both novel view synthesis and camera pose estimation.
scenes Ours COLMAP+3DGS
[…]
Church 29.39 0.92 0.10 29.93 0.93 0.09
Barn 31.98 0.94 0.08 31.08 0.95 0.07
Museum 31.92 0.94 0.08 34.47 0.96 0.05
Family 32.22 0.95 0.08 27.93 0.92 0.11
Horse 30.33 0.94 0.09 20.91 0.77 0.23
Ballroom 35.03 0.97 0.03 34.48 0.96 0.04
Francis 33.39 0.92 0.15 32.64 0.92 0.15
Ignatius 29.25 0.90 0.11 30.20 0.93 0.08
mean 31.68 0.94 0.09 30.20 0.92 0.10
Table 8: Comparison to 3DGS trained with COLMAP poses.
@@
Effectiveness of Intrinsic Optimization. Accurate camera intrinsics resolve scale ambiguity in 3DGS models, leading to improved novel view synthesis performance. As shown in Tab. 4, our method produces more accurate camera intrinsics (i.e., FoV) compared to COLMAP with original 3DGS. We also performed an ablation study with a fixed camera FoV of $60^{\circ}$ and without further optimization. The results, shown in Tab. 9, indicate a 21.4% average decrease in PSNR, due to inaccurate camera intrinsics affecting pose estimation and introducing scale ambiguity.
classroom lego_c2 livingroom bedroom mean
Ours 36.26 29.36 33.52 31.17 32.58
COLMAP+3DGS 35.81 28.77 32.74 31.73 32.26
Ours with fixed FoV (60°) 34.76 20.96 26.24 21.20 25.79
Table 9: Effectiveness of intrinsic optimization.
@@
Comparison with COLMAP-Assisted 3DGS. We compare the synthesis quality from novel views generated by our method against the original 3DGS, where the camera intrinsics and extrinsics are estimated using COLMAP on the Tanks and Temples dataset. Tab. 8 shows that our method achieves results that slightly outperform the 3DGS model trained with COLMAP-assisted poses across all scenes. Unlike the original 3DGS, which uses a fixed camera pose for training, our method seamlessly integrates 3DGS training with camera parameter estimation, allowing the two tasks to complement each other and ultimately achieve high-quality novel view synthesis. Besides, on a sequence with low-texture areas, COLMAP fails to estimate correct poses, which results in artifacts as shown in Fig. 7.
Figure 7: COLMAP failure case on Synthetic dataset.
@@
Figure 8: Visualization of ablation study on track losses in CO3D-V2. The result without track loss appears blurry in novel view synthesis.
6 Conclusion
In this work, for the first time, we integrate global tracks with 3DGS and propose a joint optimization framework for COLMAP-Free 3DGS training pipeline. Specifically, this is to avoid the instability, pose drift, and the limitations in handling complex scenes that arise from using local constraints in prior works. In our design, we introduce 2D and 3D track losses to simultaneously constrain geometric consistency in both the 2D and 3D spaces, while also constraining the scale of the track Gaussians to ensure higher spatial accuracy of 3D Gaussians. Additionally, we derive the gradient of the camera intrinsics, an important factor that was previously ignored, allowing the joint optimization of camera parameters and 3DGS to be fully differentiable. We have conducted extensive experiments on challenging datasets, where cameras have severe movements, to validate the effectiveness of our method. However, our current method does not yet support different intrinsic matrices for different views or camera distortion, which we plan to address in the future work.

### v3

@@ Introduction
Table 1: NVS results on Tanks and Temples, CO3D-V2 and Synthetic. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are gray background.
Experiments
Experimental Setup
Datasets. We conduct experiments on two real-world datasets, CO3D-V2 (co3d) and Tanks and Temples (Tanks_Temples), and a Synthetic Dataset created by ourselves. CO3D-V2 includes thousands of videos of various objects. Following CF-3DGS (CF-3DGS-2024), we select scenes with significant camera movements to demonstrate our robustness. Tanks and Temples used in CF-3DGS is overly smooth and lacks camera motions which are common in real-world scenarios. We reduce the original 20 fps to 4 fps and perform sampling along a longer camera trajectory to obtain more challenging cases with large camera motions. Synthetic Dataset comprises 4 scenes with about 150 frames each created using Blender (blender2018), showcasing complex roaming and object-centric camera motions. It’s used to assess camera parameter estimation, providing ground truth for intrinsic and extrinsic parameters. For additional details on Tanks and Temples and Synthetic Dataset, see the supplementary material.
Metrics. We use standard evaluation metrics, including PSNR, SSIM (SSIM), and LPIPS (LPIPS) to evaluate the quality of novel view synthesis (NVS). For pose estimation, we rely on the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) (barf2021; bian2022nopenerf; CF-3DGS-2024). $\mathrm{RPE}_{r}$ and $\mathrm{RPE}_{t}$ are utilized to measure the accuracy of rotation and translation. To ensure the metrics are comparable on the same scale, we align the camera poses using Umeyama’s method (Umeyama) for both estimation and evaluation. For camera focal length, we convert it to the field of view (FoV) and calculate the angular error, following (zhu2023tame).
CO3D V2 46_2587_7531 110_13051_23361 245_26182_52130 415_57112_110099
@@
Table 2: Quantitative comparison of pose accuracy on CO3D-V2. The unit of $\mathrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\mathrm{RPE}_{t}$ is scaled by 100. See the supplementary material for additional results.
Implementation Details. Our implementation is primarily based on gsplat (ye2024gsplatopensourcelibrarygaussian), an accelerated 3DGS library. We modify the CUDA operator to backpropagate gradients for camera parameters. All parameters are optimized using Adam. For initialization, we optimize the relative pose between frames, and focal length. In joint process, the 3DGS parameters, absolute poses of cameras, and focal length are optimized. As vanilla 3DGS framework demonstrates inherent limitations in simultaneous optimization of all cameras within a single step. We introduce a gradient accumulation strategy (hermans2017accumulated) that enables co-optimization of the track Gaussians and all camera parameters in a bundle adjustment manner. The camera pose is represented as a combination of an axis-angle representation $\mathrm{q}\in\mathfrak{so}(3)$ and a translation vector $\mathrm{t}\in\mathbb{R}^{3}$. During training, we will clone new Gaussians from those links to the track points and apply the same training strategy as the original 3DGS (including clone, split, and delete). The number of track Gaussians is kept constant throughout training to ensure consistency with the imposed track constraints. We set $\lambda=0.2$, $\lambda_{t2d}=0.01$, $\lambda_{t3d}=0.01$ in Eq. 10 for training.
Figure 4: Qualitative comparison for NVS and pose estimation on CO3D-V2. Benefit from the accuracy of the camera pose estimation, the rendering quality of novel view synthesis obtained by our method is higher than CF-3DGS.
Experimental Results and Analysis
Novel View Synthesis. Since the camera poses of test views are unknown, we need to estimate them for rendering. Following CF-3DGS (CF-3DGS-2024), we obtain these test-view poses by minimizing the photometric error between the synthesized images and the test views using the pre-trained 3DGS model. We apply the same procedure to all baseline methods to maintain a consistent bias for a fair comparison.
We report the results on CO3D-V2 and Tanks and Temples in Tab. 1. Our method consistently outperforms all baselines on these real-world datasets with complex camera motions. Both CF-3DGS and SFGS are less effective on these challenging datasets. The main reason is that these methods rely heavily on local constraints. When the training sequences are sparsely sampled (e.g., 20 fps $\to$ 4 fps), photometric loss and relative pose estimation are highly prone to failure and become ineffective. However, our proposed track loss effectively imposes global geometric constraints under such conditions. As illustrated in Fig. 3 and 4, the advantages of our algorithm are well demonstrated, especially with large camera motions. Due to global joint optimization, multi-view geometric consistency is better maintained in the trained 3DGS model, leading to high-quality images.
For further comparison, we evaluated our method on the Synthetic Dataset in Tab. 1, which features extremely complex camera motions. One result is shown in the bottom-right of Fig. 1. In this case, the camera not only moves in multiple circles around the object but also changes significantly in the vertical direction. Our synthesized image from the novel view remains clear and sharp, whereas the CF-3DGS result is blurry with obvious artifacts.
@@
FoV() $\mathrm{RPE}_{t}$ $\mathrm{RPE}_{r}$ ATE FoV() $\mathrm{RPE}_{t}$ $\mathrm{RPE}_{r}$ ATE FoV() $\mathrm{RPE}_{t}$ $\mathrm{RPE}_{r}$ ATE FoV() $\mathrm{RPE}_{t}$ $\mathrm{RPE}_{r}$ ATE
CF-3DGS 0.993 0.588 2.436 0.07412 0.021 1.126 4.946 0.10795 0.029 0.425 2.104 0.07653 0.042 0.366 1.103 0.06284
SFGS 0.993 0.605 2.148 0.07811 0.021 1.101 4.519 0.10904 0.029 0.400 1.817 0.07356 0.042 0.343 1.075 0.06492
COLMAP 0.993 0.004 0.018 0.00023 0.021 0.009 0.026 0.00019 0.029 0.008 0.026 0.00014 0.042 0.009 0.035 0.00023
@@
Table 3: Quantitative comparison of parameter accuracy on our Synthetic Dataset. We convert the estimated camera intrinsics focal to FoV and perform the errors of FoV with ground truth (provided by our synthetic datasets). As CF-3DGS and SFGS require the camera intrinsic parameters as fixed inputs, we set them the same as COLMAP+3DGS.
@@
Table 4: Ablation study on our Synthetic Dataset.
Scenes Ours Ours(FoV=60°) COLMAP+3DGS
PSNR LPIPS PSNR LPIPS PSNR LPIPS
classroom 36.26 0.15 34.25 0.16 35.81 0.15
lego_c2 29.36 0.12 23.82 0.26 28.77 0.15
livingroom 33.52 0.24 26.31 0.32 32.74 0.27
bedroom 31.17 0.13 24.01 0.25 31.73 0.13
Table 5: NVS results on our Synthetic Dataset compared with Fixed FoV and COLMAP-assisted 3DGS.
Camera Parameter Estimation. First, we compare the camera pose estimation with baseline methods. In the comparison, our method only assumes a constant camera focal length across frames, while others additionally input the camera focal length. The estimated camera poses are analyzed by Procrusts as in CF-3DGS and compared with the ground-truth of training views. The quantitative results of camera pose estimation on CO3D-V2 datasets are summarized in Tab. 2. The results show that our estimated camera parameters achieve the smallest error among all methods, with the Absolute Trajectory Error (ATE) being only one-tenth of that of the second-best method. This demonstrates that our algorithm excels in scenes with complex camera motions. Compared to the baselines, the global tracking information we use eliminates accumulated errors, leading to more accurate camera pose estimation. Additionally, joint optimization enhances the stability of the estimation results.
Next, we evaluate camera parameter estimation on our Synthetic Dataset. Tab. 3 shows the estimation errors from CF-3DGS, SFGS, COLMAP, and ours. Note that CF-3DGS and SFGS use the camera FoV estimated by COLMAP. We find that our estimated camera FoVs and poses are comparable to those of COLMAP, and the camera pose error is 50 times smaller than baselines. Fig. 5 visualizes our estimated poses in different stages. Thanks to the joint optimization based on global track and the back-propagation of the gradient of the camera parameters, our approach is able to combine these two tasks, reducing the input requirements.
Ablation Study
Effectiveness of Different Losses. We ablate each loss of the algorithm on Synthetic Dataset, since it has ground-truth camera parameters. Tab. 4 reports the average synthesis quality and camera parameter errors across different algorithm variants (see supplementary material for details). First, any variant of our algorithm (Variant 4, 5, 6) is better than the baseline methods (Variant 2, 3) in synthesis quality and absolute camera position. Second, Variant 4 shows that 2D track loss plays a crucial role in the entire joint optimization. When 2D track loss is not used, compared with the final method (Variant 4 vs. 6), there is a significant decrease in synthesis quality (18.18 vs. 32.58), and the camera parameter error is significantly larger (2.617 vs. 0.015). This shows that the reprojection error constrained by global consistency can significantly enhance the camera parameter estimation, thereby improving the 3DGS training effect and improving the new perspective synthesis ability, as illustrated in Fig. 6. In addition, the results of Variant 5 vs. 6 show that 3D track loss can further enhance the geometric consistency of 3DGS. Using 3D track loss, the PSNR of NVS can be further improved by 0.2 dB, and the ATE of the camera can be reduced by an order of magnitude.
@@
Effectiveness of Intrinsic Optimization. Accurate camera intrinsics resolve scale ambiguity in 3DGS models, leading to improved NVS performance. As shown in Tab. 3, our method produces more accurate intrinsics (i.e., FoV) compared to COLMAP with vanilla 3DGS. We also performed an ablation study with a fixed camera FoV of $60^{\circ}$ and without further optimization. The results, shown in Tab. 5, indicate a 21.4% average decrease in PSNR, due to scale ambiguity introduced by inaccurate camera intrinsics.
@@
Comparison with COLMAP-Assisted 3DGS. We compare the NVS results generated by our method against vanilla 3DGS, where the camera intrinsics and extrinsics are estimated using COLMAP on Synthetic Dataset. Tab. 5 shows that our method outperforms COLMAP-assisted 3DGS in most scenes. Unlike the original 3DGS, which uses a fixed camera pose for training, our method seamlessly integrates 3DGS training with camera parameter estimation, allowing the two tasks to complement each other and ultimately achieve high-quality novel view synthesis.
@@
Figure 6: Visualization of ablation study on Synthetic Dataset. Without 2D/3D track loss or intrinsic optimization, the result appears blurry in novel view synthesis.
Conclusion
We presented TrackGS, a novel framework that successfully integrates global feature tracks with 3D Gaussian Splatting to achieve COLMAP-free novel view synthesis. By establishing geometric constraints through track Gaussians and introducing novel 2D/3D consistency losses, our method jointly optimizes camera parameters and scene representation with high accuracy. The differentiable formulation of camera intrinsics completes the pipeline, removing all dependencies on precomputed parameters. Experiments demonstrate significant improvements over state-of-the-art methods in both rendering quality and pose estimation accuracy across diverse scenarios. Future extensions include extending the framework to per-view intrinsics, distortion modeling, and dynamic scene support.
Acknowledgements
This work was partly supported by the Anhui Provincial Natural Science Foundation (2508085MA001).